# Making Images Real Again: A Comprehensive Survey on Deep Image Composition

### v7

@@ VII Generative Composition
As the diffusion models [152] pretrained on large-scale dataset [155] become popular in various image generation and editing tasks, generative image composition (object compositing) has attracted growing research interest. In contrast with previous methods which perform one or multiple sub-tasks sequentially, generative image composition is a combinatorial task which performs multiple sub-tasks (e.g., image blending, image harmonization, shadow generation) parallelly through one unified model. Given a foreground, a background, and a bounding box indicating the foreground placement, generative image composition targets at directly producing a realistic composite image with the foreground naturally and harmoniously merged into the background.
@@ VII-A2 Training-based Methods
The second group of methods [222, 167, 250, 234, 230, 22, 98, 120, 178, 251, 208] require training or finetuning. They can be further divided into two subgroups, according to whether object-specific finetuning is indispensable. The first subgroup of methods [222, 167, 250, 234, 230, 22] train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. Their models need to take one or more reference images of the inserted foreground object as input. In the testing period, if a few images containing the foreground object are available, we can optionally finetune the pretrained model on these images and may achieve better performance. The second subgroup of methods [23, 120, 98, 153] do not train the model on large-scale dataset. Most of them associate the target object with one rare token, so object-specific finetuning (training on a few images of the same foreground object) is indispensable for these methods. Because the first subgroup is dominant, we will mainly introduce the first subgroup of methods.
Among the first subgroup of methods, the pioneering works like PbE [222] and ObjectStitch [167] construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [83], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. ObjectStitch [167] is extended to MureObjectStitch [20] to support multiple foreground reference images. Kulal et al. [82] adopted a similar approach, but focused on human generation. Some subsequent methods focuses on enhancing the ability of detail preservation. For example, Zhang et al. [234] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Chen et al. [22] extracted high-frequency information for better detail preservation. Yu et al. [228] designed a cycle formed by adding object and removing object, which is expected to facilitate each other. Recently, inspired by in-context learning, several recent works [191, 166, 67] explored image composition task under the framework of in-context learning. Specifically, noisy latent, background, foreground are equally fed into diffusion model and interact with each other. Different types of input are added with different levels of noise.
@@ VII-B Datasets and Evaluation Metrics
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [222, 167, 250] used FID [60] to measure the distribution discrepancy between generated images and real images. Quality scores [51, 205] are used to evaluate the authenticity of each image. CLIP score [147] or DINO score [13] are used to measure the similarity between generated foreground and reference foreground. User study should also be conducted to evaluate different aspects for the generated images.
@@ VII-C Experiments
From Fig. 18, we can see that [222] can generate foreground object with reasonable view/pose but fails to preserve the appearance details. Compared with [22, 234, 20], [166] can preserve very small details (e.g., the letters on the cake in row 3, the pattern on the waste container in row 5), probably due to the detail preservation ability of in-context learning. The methods [22, 234, 166] exhibit notable copy-and-paste effect and heavily rely on suitable foreground reference image. If the view/pose of foreground reference image is not compatible with the background, these methods [22, 234, 166] are weak in adjusting the foreground view/pose to match the background.
Additionally, these methods require manual specification of reasonable bounding box and most of them cannot generate high-quality shadow/reflection that falls out of the bounding box, so object placement and shadow/reflection generation are actually not fully accomplished.
@@ VIII Foreground Object Search
Figure 19: Illustration of foreground object search. Given a background image with query bounding box (yellow), foreground object search aims to find compatible foreground objects of a specified category from a library, which is composited with the background to produce a realistic composite image.
@@
The goal of foreground object search (FOS) is retrieving suitable foregrounds from a foreground library, which are compatible with the background in terms of illumination, geometry, and semantics. The FOS task is illustrated in Fig. 19. Finding compatible foregrounds can greatly alleviate the burden of creating realistic composite images, which is complementary with the other image composition techniques. FOS task can be divided into constrained or unconstrained according to whether specifying the foreground category.
@@ VIII-A Traditional Methods
Early works [86, 21] attempted to match each foreground with the background using hand-crafted features, but their performance is limited by the representation ability of hand-crafted features. Specifically, Lalonde et al. [86] estimated the object information (e.g., size, orientation, lighting condition) and designed matching criteria to rank all the objects in the library. Chen et al. [21] exploited the contour consistency and content consistency between foreground and background based on hand-crafted features.
@@ VIII-B Deep Learning Methods
Recent work used deep learning features for foreground retrieval. For example, Tan et al. [172] utilized deep features to capture local context particularly for person compositing. Zhu et al. [260] trained a composite image discriminator to predict the realism of composites by compositing each foreground with the background. This method is effective by using the realism of composite image to measure the foreground-background compatibility, but computing the realism of all composite images is very expensive. More recent methods [253, 254, 261, 213, 90, 236] typically trained two encoders to extract foreground feature and background feature. Then, the foreground-background compatibility is measured by calculating the distance between foreground feature and background feature. They share the similar framework, despite the difference in data preparation, network structure, and loss design. Zhang et al. [236] observed that a composite image discriminator [260] can perform much better than two encoders, so they developed a teacher-student network which distills composite image feature from discriminator to the interaction output of foreground feature and background feature.
As introduced in Section I, the foreground and background in a composite image have multiple types of inconsistencies. The existing FOS works considered different sets of inconsistencies between background and foreground. For example, the methods [253, 254] considered the semantic consistency. The methods  [90, 236] considered the geometry consistency and semantic consistency. Besides the geometry and semantic consistency, some other methods [213, 261] additionally considered style consistency [213] or lighting consistency [261].
@@ VIII-C Datasets and Evaluation Metrics
Early FOS works [253, 254, 213, 261] did not release their datasets. Zhang et al. [236] contributed two datasets: S-FOSD and R-FOSD, which contain synthetic composite images and real composite images respectively. In S-FOSD dataset, Zhang et al. [236] segment one foreground object from a real image and fill its bounding box with image mean values to get the background. For each background image, the foreground object from the same image is deemed as ground-truth. In R-FOSD dataset, Zhang et al. [236] collect images from Internet as background images and draw a bounding box at the expected foreground location as query bounding box. R-FOSD dataset shares the same foregrounds with the test set of S-FOSD dataset. Zhang et al. [236] employ human annotators to label the compatibility of each pair of background and foreground. In comparison, S-FOSD dataset is low-cost and highly scalable, but has neither complete background nor ground-truth negative samples. R-FOSD dataset has complete background image with accurately annotated positive and negative foregrounds, but is unscalable due to the high annotation cost.
On synthetic composite image dataset, Recall@k (R@k) is adopted as evaluation metric [253, 261, 236], which represents the percentage of background queries whose ground-truth foreground appears in top $k$ retrievals. On real composite image dataset, mean Average Precision (mAP), mAP@20, and Precision@k (P@k) are adopted as evaluation metrics [254, 261, 236].
@@ VIII-D Experiments
We evaluate different methods on S-FOSD dataset and R-FOSD dataset [236]. Specifically, we train on S-FOSD training set, while testing on S-FOSD test set and R-FOSD test set. The retrieval results of CFO [253], UFO [254], GALA [261], FFR [213], and DiscoFOS [236] are shown in Fig. 20. The results show that DiscoFOS can retrieve more foregrounds that are compatible with the background.
@@ IX Conclusion
In this paper, we have conducted a comprehensive survey on image composition, which involves a variety of techniques to produce a realistic composite image. We have introduced object placement, image blending, image harmonization, shadow generation, generative image composition, foreground object search in detail. In the future, we will extend this survey to more general composition in other fields such as video composition and 3D composition.

### v8

@@ VII Generative Composition
As the diffusion models [160] pretrained on large-scale dataset [163] become popular in various image generation and editing tasks, generative image composition (object compositing) has attracted growing research interest. In contrast with previous methods which perform one or multiple sub-tasks sequentially, generative image composition is a combinatorial task which performs multiple sub-tasks (e.g., image blending, image harmonization, shadow generation) in parallel through one unified model. Given a foreground, a background, and a bounding box indicating the foreground placement, generative image composition aims to directly produce a realistic composite image with the foreground naturally and harmoniously merged into the background.
@@ VII-A2 Training-based Methods
The second group of methods [230, 175, 260, 243, 21, 239, 24, 104, 127, 186, 262, 216, 254, 231, 109] require training or finetuning. They can be further divided into two subgroups, according to whether object-specific finetuning is indispensable. The first subgroup of methods [230, 175, 260, 243, 239, 24] train a diffusion model on abundant pairs of foregrounds and backgrounds, so that it can be directly applied to a new pair of foreground and background at test time. These models need to take one or more reference images of the inserted foreground object as input. In the testing period, if a few images containing the foreground object are available, we can optionally finetune the pretrained model on these images and may achieve better performance. The second subgroup of methods [25, 127, 104, 161] do not train the model on large-scale dataset. Most of them associate the target object with one rare token, so object-specific finetuning (training on a few images of the same foreground object) is indispensable for these methods. Because the first subgroup is dominant, we will mainly introduce the first subgroup of methods.
Among the first subgroup of methods, the pioneering works like PbE [230] and ObjectStitch [175] construct massive training triplets of foregrounds, backgrounds, and ground-truth real images based on large-scale image datasets [88], in which the foregrounds are cropped from real images followed by color and geometry perturbation. Then, they adapt conditional diffusion model to this task. In particular, the background image, bounding box mask, and noisy image are concatenated as input, while the foreground is injected into the network via cross-attention. ObjectStitch [175] is extended to MureObjectStitch [20] to support multiple foreground reference images. Kulal et al. [87] adopted a similar approach, but focused on human generation. Some subsequent methods focus on enhancing the ability of detail preservation. For example, Zhang et al. [243] proposed global-and-local fusion, in which shallow foreground features are used to enhance the details. Chen et al. [24] extracted high-frequency information for better detail preservation. Yu et al. [237] designed a cycle formed by adding object and removing object, which is expected to facilitate each other. Recently, inspired by in-context learning, several recent works [199, 174, 72] explored image composition task under the framework of in-context learning. Specifically, noisy latent, background, foreground are equally fed into diffusion model and interact with each other. Different types of input are added with different levels of noise.
@@ VII-B Datasets and Evaluation Metrics
In real-world application scenarios, there exist no ground-truth images for a pair of foreground and background, so we cannot calculate the distance between generated image and ground-truth image. Therefore, previous works [230, 175, 260] used FID [65] to measure the distribution discrepancy between generated images and real images. Quality metrics [56, 213] are used to evaluate the authenticity of each image. CLIP score [155] or DINO score [13] are used to measure the similarity between generated foreground and reference foreground. User study should also be conducted to evaluate different aspects for the generated images.
@@ VII-C Experiments
From Fig. 18, we can see that [230] can generate foreground object with reasonable view/pose but fails to preserve the appearance details. Compared with [24, 243, 20], [174] can preserve very small details (e.g., the letters on the cake in row 3, the pattern on the waste container in row 5), probably due to the detail preservation ability of in-context learning. The methods [24, 243] exhibit notable copy-and-paste effect and heavily rely on suitable foreground reference image. If the view/pose of foreground reference image is not compatible with the background, these methods [24, 243] are weak in adjusting the foreground view/pose to match the background. Compared with [24, 243] , [174] has better ability to adjust the foreground view/pose, but may still fail in the challenging cases.
For practical usage, in-context learning methods are recommended to preserve the subtle foreground details. When multiple foreground reference images are available, it is suggested to select the one relatively compatible with the background in advance, which could significantly increase the success rate of generative image composition.
@@ VIII Foreground Object Search
Figure 19: Illustration of foreground object search. Given a background image with query bounding box (yellow), foreground object search aims to find compatible foreground objects of a specified category from a library, which is combined with the background to produce a realistic composite image.
@@
The goal of foreground object search (FOS) is to retrieve suitable foreground objects from a foreground library, which are compatible with the background in terms of illumination, geometry, and semantics. The FOS task is illustrated in Fig. 19. Finding compatible foreground objects can significantly reduce the effort required to create realistic composite images, complementing other image composition techniques. FOS task can be divided into constrained or unconstrained depending on whether the foreground category is specified.
@@ VIII-A Traditional Methods
Early works [91, 23] attempted to match each foreground with the background using hand-crafted features, but their performance is limited by the expressiveness of hand-crafted features. Specifically, Lalonde et al. [91] estimated the object information (e.g., size, orientation, lighting condition) and designed matching criteria to rank all the objects in the library. Chen et al. [23] exploited the contour consistency and content consistency between foreground and background based on hand-crafted features.
@@ VIII-B Deep Learning Methods
Recent work used deep learning features for foreground retrieval. For example, Tan et al. [180] utilized deep features to capture local context particularly for person compositing. Zhu et al. [271] trained a composite image discriminator to predict the realism of composites by compositing each foreground with the background. This method is effective in using the realism of composite images to measure the foreground-background compatibility, but computing the realism of all composite images is very expensive. More recent methods [264, 265, 272, 221, 95, 245] typically trained two encoders to extract foreground feature and background feature. Then, the foreground-background compatibility is measured by calculating the distance between foreground feature and background feature. They share a similar framework, despite the difference in data preparation, network structure, and loss design. Zhang et al. [245] observed that a composite image discriminator [271] can perform much better than two encoders, so they developed a teacher-student network which distills composite image feature from the discriminator to the interaction output of foreground feature and background feature.
As introduced in Section I, the foreground and background in a composite image have multiple types of inconsistencies. The existing FOS works considered different sets of inconsistencies between background and foreground. For example, the methods [264, 265] considered the semantic consistency. The methods  [95, 245] considered the geometric consistency and semantic consistency. Besides the geometric and semantic consistency, some other methods [221, 272] additionally considered style consistency [221] or illumination consistency [272].
@@ VIII-C Datasets and Evaluation Metrics
Early FOS studies [264, 265, 221, 272] did not release their datasets. Zhang et al. [245] contributed two datasets: S-FOSD and R-FOSD, which contain synthetic composite images and real composite images respectively. In S-FOSD dataset, Zhang et al. [245] segment one foreground object from a real image and fill its bounding box with image mean values to get the background. For each background image, the foreground object from the same image is deemed as ground-truth. In R-FOSD dataset, Zhang et al. [245] collect images from Internet as background images and draw a bounding box at the expected foreground location as query bounding box. R-FOSD dataset uses the same foregrounds with the test set of S-FOSD dataset. Zhang et al. [245] employ human annotators to label the compatibility of each pair of background and foreground. In comparison, S-FOSD dataset is low-cost and highly scalable, but does not contain complete background images or ground-truth negative samples. R-FOSD dataset has complete background image with accurately annotated positive and negative foregrounds, but is unscalable due to the high annotation cost.
On the synthetic composite image dataset, Recall@k (R@k) is adopted as an evaluation metric [264, 272, 245], which represents the percentage of background queries whose ground-truth foreground appears in top $k$ retrievals. On the real composite image dataset, mean Average Precision (mAP), mAP@20, and Precision@k (P@k) are adopted as an evaluation metrics [265, 272, 245].
@@ VIII-D Experiments
We evaluate different methods on S-FOSD dataset and R-FOSD dataset [245]. Specifically, we train on S-FOSD training set, while testing on S-FOSD and R-FOSD test sets. The retrieval results of CFO [264], UFO [265], GALA [272], FFR [221], and DiscoFOS [245] are shown in Fig. 20. In each example, we show the background image with query bounding box (gray square or yellow box) on the left side and top five retrieved results of different methods on the right side. The results show that DiscoFOS can retrieve more foregrounds that are compatible with the background geometrically and semantically.
@@ IX Conclusion
In this paper, we have conducted a comprehensive survey on image composition, which involves a variety of techniques to produce a realistic composite image. We have introduced object placement, image blending, image harmonization, shadow generation, generative image composition, and foreground object search. In the future, we will extend this survey to broader composition tasks in related fields, such as video composition and 3D composition.